### v2

# Open-World Task and Motion Planning via Vision-Language Model Inferred Constraints

Abstract
Foundation models trained on internet-scale data, such as Vision-Language Models (VLMs), excel at performing a wide variety of common sense tasks like visual question answering. Despite their impressive capabilities, these models cannot currently be directly applied to challenging robot manipulation problems that require complex and precise continuous reasoning over long horizons. Task and Motion Planning (TAMP) systems can control high-dimensional continuous systems over long horizons via a hybrid search over traditional primitive robot skills. However, these systems require detailed models of how the robot can impact its environment, preventing them from directly interpreting and addressing novel human objectives, for example, an arbitrary natural language goal. We propose deploying VLMs within TAMP systems by having them generate discrete and continuous language-parameterized constraints that enable TAMP to reason about open-world concepts. Specifically, we propose algorithms for VLM partial planning that constrain a TAMP system’s discrete temporal search and VLM continuous constraints interpretation to augment the traditional manipulation constraints that TAMP systems seek to satisfy. Experiments demonstrate that our approach — OWL-TAMP — outperforms several related baselines, including those that solely use TAMP or VLMs for planning, across several long-horizon manipulation tasks specified directly through natural language. We additionally demonstrate that our approach is compatible with a variety of TAMP systems and can be deployed to solve challenging manipulation tasks on real-world hardware. Project website: https://owl-tamp.csail.mit.edu/. Work partially conducted at NVIDIA Research. Correspondence to njk@csail.mit.edu, cgarrett@nvidia.com.
I Introduction
The advent of foundation models trained on internet-scale data has led to unprecedented progress on traditionally-hard tasks in vision and natural language. Current Large Language Models (LLMs) and Vision-Language Models (VLMs) are able to complete text from partial specifications, answer questions about images, and even solve challenging word problems that require reasoning and common sense [1, 2, 3]. This impressive performance has inspired several systems that attempt to use existing pretrained models in robotics [4, 5, 6, 7]. Such systems exhibit impressive flexibility: unlike classical robotics approaches, they are able to accomplish novel goals specified by natural language or images. However, currently no publicly-available foundation models exist that can directly output continuous values (e.g. joint angles, grasps, placements), which are critical for interacting with the physical world.
In contrast, classical Task and Motion Planning (TAMP) systems are capable of solving complex and long-horizon tasks ranging from setting a dining table to 3D printing complex structures [8, 9, 10]. These systems leverage models of the robot and its environment to explicitly reason about both discrete and continuous values in robotics problems. While such systems are powerful on the set of problems they have been designed for, they do not transfer to novel problems for which their models are unspecified. Enabling a TAMP system to solve novel problems often requires manually extending the underlying model, which is tedious and not scalable when operating in unstructured human environments.
We are interested in combining the complementary benefits of foundation-models and TAMP to tackle long-horizon manipulation tasks that are open world, namely where the vocabulary of objectives is unbounded. Specifically, we assume tasks are specified in natural language, which may involve concepts that an underlying TAMP system does not have built-in, but which can be achieved by chaining together robot motion primitives the TAMP system possesses. As an example, a TAMP system that is capable of accomplishing pick-and-place tasks expects goals in the form of logical expressions involving predicates like On(apple, plate). Consider the goal in Figure LABEL:fig:system: “Put the orange on the table where the apple initially is”. This goal cannot be expressed in terms of On, and thus there would be no way a TAMP system could solve it, even though it could be accomplished by a sequence of pick-place primitives. A pure VLM system would also struggle with this task since it must not only predict that the apple needs to be moved out of the way before the orange can be placed, but also continuous robot motions that realize this.
Our key insight is that we can integrate the discrete-continuous planning of TAMP systems and common sense reasoning of VLMs through the contract of constraints. In particular, VLMs are capable of mapping a very wide range of open world expressions into discrete action sequences (e.g. that a potato must be cooked before it can be served), and code that represents continuous constraints over important decision variables (e.g. valid poses of the egg such that it is inside an oven). These constraints can be readily integrated with existing constraints (e.g. avoiding collisions, respecting kinematics) within off-the-shelf TAMP systems. Thus, the overall system is able to generate solutions that not only respect constraints derived from the open world goal, but also are physically feasible on robot hardware.
We propose OWL-TAMP (Open-World Language-based TAMP), an approach that integrates open world concepts via constraint generation into a TAMP system with traditional robotics operations and constraints. Our key contributions towards this framework are: (1) a method for generating constraints on action sequences to specify partial plans with language descriptions; (2) a method for generating constraints on continuous variables affected within the partial plan from (1); and (3) combining both (1) and (2) within a TAMP system. We test our framework on a variety of open world tasks in simulation and find that it is able to solve open world tasks with a higher success rate than several ablations, including a pure VLM or pure TAMP system respectively. We also demonstrate that OWL-TAMP enables a real-world robot to solve complex, long-horizon manipulation tasks specified through language directly from sensor input.
II Related Work
Task and Motion Planning (TAMP). A standard approach to solving long-horizon robot manipulation tasks via joint discrete-continuous reasoning is TAMP [11, 12]. Task planners can be made to solve open-world tasks expressed in natural language by translating the language into a symbolic goal expression conducive to planning [13]. However, this approach relies on the assumptions that: (1) the goal expression can be translated exactly into an expression over a set of predicates and (2) the translated expression can be achieved by planning with specific symbolic operators. Enabling TAMP systems to solve a wider range of tasks requires expanding their symbolic vocabulary either manually, or via learning [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. By contrast, our approach leverages VLMs to directly write constraint functions that expand the set of tasks a TAMP system can solve without directly expanding its base symbolic vocabulary.
LLMs and VLMs for Robotic Manipulation. Recent work has leveraged internet-scale foundation models to enable complex short-horizon visuomotor behaviors [26, 27, 4, 28, 29]. These approaches seek to solve relatively short-horizon tasks (such as pouring water into a cup or folding a t-shirt) whereas we are interested in longer horizon tasks composed of multiple atomic behaviors. Additionally, to the extent that these approaches do handle multi-step tasks (e.g. pick-and-place tasks), they often make assumptions or leverage heuristics specific to the particular class of tasks (e.g. that the task always consists of a series of picks followed by placements) [4], or require learning in simulation, and are thus not zero-shot [28, 29]. By contrast, our work proposes a clear contract for combining VLMs with generic domain-independent TAMP systems, with no requirement for learning in simulation, and minimal assumptions on the class of tasks that the system may be asked to perform.
LLMs and VLMs for Robot Task Planning. Another line of work has leveraged foundation models to solve long horizon planning problems. These works have focused on the problem of sequencing discrete skills in restricted settings where continuous parameters are either unnecessary, or can be easily determined by simple heuristics [5, 30, 31, 6, 32]. Some works have considered tasks where continuous parameters are important [32, 33, 34, 35, 36, 37, 38], but have relied on foundation models to either directly output such parameters, or to write code to do so. Thus, these approaches have generally only been able to handle problems in which the continuous parameters can be easily determined, most do not consider resolving complex constraints like reachability, kinematic feasibility or grasp stability. The few approaches that solve complex tasks that require precise discrete and continuous parameters [39, 40, 41, 42, 7] either have limited ability to handle cases where the system must consider alternative symbolic plans to achieve a goal (e.g. moving an object out of the way in order to place an object), or generally struggle with long-horizon tasks (where explicit search tends to perform better than LLMs [43, 44]). By contrast, our approach leverages off-the-shelf TAMP systems for planning and is thus able to solve longer horizon tasks with more complex constraints. Concurrent work [45] attempts to define a similar constraint-based contract between LLMs and TAMP systems for manipulation tasks, but focuses only on discrete constraints whereas our approach is able to leverage VLMs for both discrete and continuous constraint generation.
III Problem Setup
We adopt a model-based mixed discrete-continuous planning approach to controlling a robot to solve open-world tasks. We assume a planning model which contains commonplace manipulation primitives applicable across a very wide range of tasks and leverage foundation models to extend the core model to reason about novel, task-specific dynamics and constraints. Intuitively, the underlying planning model should capture generic dynamics and constraints (e.g. inverse kinematics and reachability, collision constraints) that apply across any task a robot might be faced with, while the foundation model should provide additional task-specific constraints (e.g. that an object must be placed in a pan for it to be ‘cooked’, that serving coffee in a mug requires that mug be upright) that serve to specialize the planning model to the given situation.
We model the system using a Planning Domain Definition Language (PDDL)-style [46] factored action language, which represents states and actions in terms of predicates. Our approach is not limited to this representational choice; it is compatible with multiple planning frameworks, e.g. PDDLStream [9] and SeSaME [12, 8, 16]. In PDDL, state variables are represented as literals, true or false evaluations of predicates for particular values of their parameters.
We use a single robot acting in a simplified manipulation domain as a pedagogical running example. Because robotics inherently involves continuous values, we consider not only discrete parameter types but also continuous ones, namely: obj - a discrete manipulable object $o$, conf - a continuous robot configuration $q\in\mathbf{R}^{d}$, traj - a continuous robot trajectory comprised of a sequence $n$ of configurations $\tau\in\mathbf{R}^{nd}$, grasp - a continuous object grasp pose $g\in\mathrm{SE}(3)$, and pose - a continuous object placement pose $p\in\mathrm{SE}(3)$.
The fluent predicates, i.e. predicates with truth values that can change over time, are: AtConf($q$: conf) - the robot is currently at configuration $q$, HandEmpty() - the robot’s hand is currently empty, AtPose($o$: obj, $p$: pose) - object $o$ is currently at placement pose $p$, and AtGrasp($o$: obj, $g$: grasp) - object $o$ is currently grasped with grasp pose $g$. From these predicates, we can describe states, which are represented by true literals. For example, the initial state in a domain with a single object apple might be: $s_{0}=[{\texttt{AtConf}}(q_{0}),{\texttt{HandEmpty}}(),{\texttt{AtPose}}({\texttt{apple}},p_{0}),...].$
Figure 1: Hybrid traditional and VLM constraint satisfaction. The constraint network for a candidate plan to pick and place the apple that achieves the goal “put the orange on the table where the apple initially is” in Figure LABEL:fig:system. Due to the Collision and VLMPose constraints, this constraint network isn’t satisfiable; however, a plan that first moves apple out of the way before placing apple is.
Parameterized actions, which the robot can apply to affect a change in a state, are defined by a name, list of typed parameters, list of static literal constraints (con) that the parameters must satisfy, list of fluent literal preconditions (pre) that must hold before applying the action, and list of fluent literal effects (eff) that hold in the state after applying the action. The actions move and attach model the robot moving between two configurations and attaching an object to itself, for example, by grasping it.
move$(q_{1}:\text{conf},\;q_{2}:\text{conf},\;\tau:\text{traj})$
con: [Motion$(q_{1},\tau,q_{2})$]
pre: [AtConf$(q_{1})$]
eff: [AtConf$(q_{2})$, $\neg$¬AtConf$(q_{1})$]
attach$(o:\text{obj},\;p:\text{pose},\;g:\text{grasp},\;q:\text{conf})$
con: [Kin$(q,o,g,p)$]
pre: [AtPose$(o,p)$, HandEmpty()()( ), AtConf$(q)$]
eff: [AtGrasp$(o,g)$, $\neg$¬AtPose$(o,p)$, $\neg$¬HandEmpty()()( )]
Ground action instances of these parameterized actions must satisfy the following static predicates: Motion($q_{1}$: conf, $\tau$: traj, $q_{2}$: conf) - $\tau$ is a valid trajectory that connects configurations $q_{1}$ and $q_{2}$, and Kin($q$: conf, $o$: obj, $g$: grasp, $p$: placement) - configuration $q$ satisfies a kinematics constraint with placement pose $p$ when object $o$ is grasped with grasp pose $g$.
III-A Open World Predicates and Actions
So far, we’ve described a small and finite set of traditional TAMP predicates and actions. Intuitively, these correspond to generic dynamics and constraints that a robot encounters due to its embodiment in the physical world. However, we are interested in modeling and planning with open-world concepts that are environment or task specific. To do so, we propose parameterizing select predicates and actions with an additional type, a description $d$. Descriptions modify the semantics of predicates and actions to respect an open-world natural-language instruction. Descriptions help specialize the overly general robot interactions (e.g. moving without collision, grasping stably) in the traditional planning model to achieve novel outcomes. Overall, this strategy can be seen as bootstrapping an unbounded set of predicates and actions from a finite set by leveraging language itself as a parameter.
Consider the VLMPose($d$: description, $o$: obj, $p$: pose) constraint, which is true if object $o$ at placement $p$ satisfies description $d$. Some example descriptions $d$ are: “orange at the center of the table”, “orange at the apple’s initial location”, and “orange as far way from the robot as possible”. Using this constraint, we can formulate a detach action, which involves the robot releasing object $o$ according to the description $d$. This can correspond to placing the object on a surface, stacking the object on another object, dropping the object in a bin, inserting the object into an outlet, etc.
detach$(d:\text{description},\;o:\text{obj},\;g:\text{grasp},\;p:\text{pose},\;q:\text{conf})$
con: [Kin$(q,o,g,p)$, VLMPose$(d,o,p)$]
pre: [AtPose$(o,p)$, HandEmpty()()( ), AtConf$(q)$,
$\neg\exists o^{\prime},p^{\prime}.\;{\texttt{AtPose}}(o^{\prime},p^{\prime})\wedge{\texttt{Collision}}(o,p,o^{\prime},p^{\prime})$]
eff: [AtGrasp$(o,g)$, $\neg$¬AtPose$(o,p)$, $\neg$¬HandEmpty()()( )]
We can define additional parameterized actions that model different interaction types, such as an action that moves a cup through waypoints to fill it up or pour out of it  [^1] We use demonstrate this in the ‘Mug3’ and ‘SoupPour’ tasks in Section V..
A key advantage of our approach is that we plan with both traditional robot constraints as well as task-specific open-world constraints. Consider the problem in Figure LABEL:fig:system, where the the goal is to ‘put the orange on the table where the apple initially is”. Figure 1 (left) displays the simplified constraint network [11], a bipartite graph from free action parameters (in bold) to the action constraints they are involved in (conf), induced by a plan that directly picks and places the apple:
$\pi$ $=[...,{\texttt{attach}}({\texttt{apple}},p_{0}^{A},\bm{g},\bm{q_{1}}),...,{\texttt{detach}}($
[…]
This constraint network is unsatisfiable because the VLMPose constraint restricts the set of placements that satisfy the task and the Collision constraint prevents unsafe placements. But through the use of TAMP, our approach can backtrack over candidate plans that first move the apple to eventually find a satisfiable constraint network and ultimately a solution.
IV TAMP with Open World Concepts
We address TAMP problems $\langle s_{0},{\cal A},g\rangle$ described by an initial state $s_{0}$, set of parameterized actions ${\cal A}$, and goal $g$. Unlike traditional TAMP problems, the goal $g$ is not a logical formula over literals but rather is a goal description provided in English text. Thus, solving such problems requires translating $g$ into some form that can be used within TAMP.
One approach to this translation would be to directly prompt a VLM to output some logical formula over literals (which we will denote as $G$) from the goal description $g$ [13]. Given this, one could simply call an off-the-shelf TAMP system to achieve $G$. While this approach is straightforward, and powerful, it is limited in the kinds of tasks it is able to express in at least two ways: (1) it can only define a goal state to achieve and cannot specify intermediate behaviors or states that need to occur before the goal, and (2) it can only express goals in terms of predicates that are already built into the TAMP system.
[…]
We seek to address these limitations in the expressivity of direct translation by instead translating $g$ into more flexible discrete and continuous constraints (as depicted in Figure LABEL:fig:system). Specifically, we first prompt a VLM to supply a set of discrete constraints over open world action orderings, and then induce continuous constraints in the form of code for particular predicates (such as VLMPose) that appear in the effects or constraints of action definitions used as part of our first stage. We then incorporate these constraints into a TAMP system such that it only yields plans that satisfy these constraints. Intuitively, these constraints will be task specific and enable the system to achieve tasks it otherwise could not. Conversely, through using a TAMP system, OWL-TAMP inherits theoretical guarantees with respect to the non-VLM constraints such as plan soundness, which is critical for safety, and probabilistic completeness [47]. In the cooking task mentioned above, generating a discrete constraint that any valid plan should execute a detach(strawberry, pan) action before a detach(strawberry, bowl) action would be sufficient to enable the TAMP system to solve the task. Similarly, in the fruit sorting task, all that is required is a continuous constraint on the outcome of every detach(fruit) for a TAMP system to accomplish the underlying goal.
In what follows, we first describe our procedure for discrete constraint generation before describing how we generate continuous constraints given initial discrete constraints.
IV-A Generating discrete planning constraints with a VLM
Given a goal description $g$, we first prompt a VLM to generate a partial plan that serves as a discrete constraint on the space of TAMP solutions. To enable this, we associate a natural language description of each available action with that particular action. Although we could directly prompt a VLM for relevant actions and goals, without a list of candidates, the VLM is likely to be syntactically and semantically inaccurate. Instead, we propose to first ground the set of reachable actions $A$ and literals $L$ available to the TAMP system before prompting the VLM to return values in these sets. We use relaxed planning [48, 49] from the initial state $s_{0}$ to simultaneously ground [50, 51, 52] and explore the sets of reachable actions $A$ and literals $L$. When instantiating continuous parameters, we use placeholder values, such as optimistic values [53, 47, 9, 54], to ensure a finite set of actions are instantiated. Similarly, we use placeholders for description parameters.
Algorithm 1 VLM Task Reasoning
1:procedure vlm-task-reasoning($s_{0},{\cal A},g$)
2:    $A\leftarrow\textsc{ground-actions}(s_{0},{\cal A})$
[…]
4:    $[a_{1},...,a_{n},l_{1},..,l_{m}]\leftarrow\textsc{query-vlm}(\text{``What partial plan}\newline \text{using actions }\{A\}\text{ for goal literals }\{L\}\text{ achieves goal }\{g\}\text{?''})$” )
5:    for $i\in[1,n-1]$ do
[…]
9:    $G\leftarrow\{l_{1},..,l_{m}\}$
[…]
Algorithm 1 presents the VLM partial plan generation pseudocode. It takes in a TAMP problem $\langle s_{0},{\cal A},g\rangle$, where $g$ is a text goal description. It first grounds the set of actions $A$ reachable from $s_{0}$ using ground-actions. Then, it accumulates the set of reachable literals $L$ by taking the effects of all actions $A$. These sets can be filtered by action or predicate type if it is desired to focus VLM assistance on specific aspects of the planning problem. Then, it prompts query-vlm for a partial plan $[a_{1},...,a_{n},l_{1},...,l_{k}]$ using actions $a_{i}\in A$ and goal literals $l_{j}\in L_{m}$ that achieve the goal description $g$. Importantly, we have the VLM fill in the description parameter $d$ for each of these actions. We then transform the original TAMP problem to force solutions to admit the partial plan as a subsequence. Specifically, we create a predicate Executed that models whether the $i$th action in the plan was executed and add Executed to the effects of action $a_{i}$ and the preconditions of action $a_{i+1}$. Finally, we make the planning goal be $G=\{l_{i},...,l_{m}\}\subseteq L$ and $\textsc{Executed}(n)$, which indicates that all actions have been executed and solve the transformed TAMP problem with a generic TAMP algorithm.
Consider the cooking problem mentioned earlier in the beginning of Section IV $g=$“Cook the strawberry by putting it in the pan, then finally serve it in the bowl”. Suppose the VLM returns no goal literals, but just the partial plan:
[…]
Although the VLM plan $\vec{\pi}$ does capture the intent of the task (i.e., to place the apple in the pan before serving it), this plan is not legal because objects must be picked with the attach action before they can be detached. Fortunately, the underlying TAMP system models this, and thus providing this partial plan, along with the generated Executed predicates, to the TAMP system will result in the TAMP system generating legal plans that are at least 4 actions long.
Figure 2: Simulated RAVENS-YCB Manipulation Tasks. Intermediate states during execution of our approach on 3 simulated tasks.
Figure 3: Real-World Tasks. Two tasks from our real-world system deployment (Section V) and their corresponding goal descriptions.
IV-B Grounding Continuous Constraints with a VLM
The procedure in Section IV-A enables us to generate actions with language parameters (Section III-A) fully specified. However, in order to correctly apply these actions, we need to be able to interpret how the language parameter should affect legal action parameter values. More specifically, we need to provide an implementation for any constraint fluents (such as the VLMPose$(d,o,p)$ fluent introduced in Section III-A) that use the language description $d$.
More specifically, consider the coffee task (i.e. where $g=$ “Can you setup the cup on the table so I can properly pour coffee into it?”), and suppose the discrete generation procedure has produced a plan that contains the following action:
[…]
ensuring it is upright and positioned
[…]
To properly implement this action, we must ensure that the placement pose $p$ of detach action obeys the description $d$ of being “stably on the table and upright”. To this end, we prompt a VLM to generate code to implement a test on the pose $p$ directly that outputs a Boolean value (and can thus be used as part of VLMPose):
1def test_poses(p) -> bool:
2 ontop_table_bounds = modify_pose_bounds_to_be_ontop _of_object(’mug’, ’table’)
3 mug_on_table = position_within_bounds(mug.pose, ontop_table_bounds)
4 upright_orientation = abs(mug.pose.roll) < 0.1 and abs(mug.pose.pitch) < 0.1
5 return mug_on_table and upright_orientation
Given such a function, we can implement the VLMPose($d,o,p$) predicate by simply calling this function and passing in the pose pose $p$ at which the mug object is being placed. The description $d$ is passed into the VLM to generate this function. Given this implementation on VLMPose, the TAMP system will be constrained to solutions that respect this continuous constraint, in line with the intent of the task. Although we focus on Boolean functions as action constraints, our approach can be straightforwardly applied to nonnegative functions as action costs to, for example, minimize the distance from a placement to a table edge.
We additionally found that it is helpful to have the VLM output continuous constraints corresponding to the goal description $g$ itself, and then use these to output constraints on each of the discrete actions. We then feed its output from this step as part of the prompts for it to output constraints on every other action with description $d$ and a constraint fluent requiring a VLM implementation. Additional details related to prompting and constraint implementation are described in Appendix A-F, and a full walkthrough of generating constraints on an example task is shown in Appendix A-B2.
V Experiments and System Demonstration
Our experiments measure how well our approach is able to solve tasks with natural-language goals and requiring satisfying a wide range of different constraints. We compare our approach quantitatively against several ablations and baselines in simulation, and then implement and demonstrate its capabilities on real-world hardware.
Tasks
Method Berry1 Citrus Berry2 BerryCook FruitSort Coffee Mug1 Mug2 Mug3 SoupPour Overall
CaP 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
CaP-sample 100% 20% 20% 0% 0% 0% 0% 0% 0% 0% 14%
No VLM 100% 100% 100% 0% 0% 0% 0% 0% 20% 0% 32%
No sample 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
No disc. 100% 100% 100% 0% 0% 0% 0% 0% 20% 0% 32%
No cont. 100% 100% 100% 100% 10% 60% 70% 0% 0% 20% 56%
No back. 100% 90% 100% 60% 80% 100% 40% 30% 0% 0% 60%
OWL-TAMP 100% 100% 100% 60% 100% 100% 100% 100% 70% 90% 92%
TABLE I: Success rates on all tasks. We report the success rate (i.e., whether the task was successfully accomplished, as measured by manually-written checks), across 10 random seeds. We run a one-tailed z-test with $\alpha=0.1$ and bold all top-scoring success rates. The final column displays the average success rate across all trials on all tasks.
Tasks. We introduce the tasks we tested our approach and baselines on, with additional details and discussion in Appendix A. These tasks are all implemented within the RAVENS-YCB Manipulation environment adapted from previous work [7, 55]. The environment features a 6 DoF UR5 robot arm with a Robotiq 2F-85 gripper in front of a tabletop strewn with objects from the YCB dataset [56], as visualized in Figure 3. The $10$ tasks are listed here in approximate order of difficulty, where more difficult tasks involve satisfying a larger number and complexity of constraints.
  • $\bullet$
    Berry1: “put the strawberry onto the light-grey region at the center of the table”. The optimal solution to this task requires a sequence of 2 skills.
  • $\bullet$
    Citrus: “pack the citrus fruit onto the plate”. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    Berry2: Same as Berry1, but requires moving an obstacle out of the way of the light-grey region. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    BerryCook: “Cook the strawberry by putting it in the pan, then finally simply place it in the bowl. The strawberry should only be in the bowl at the end!”. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    FruitSort: “Put all the fruit to the left of the line bisecting the table”. The optimal solution to this task requires a sequence of 6 skills.
  • $\bullet$
    Coffee: “I want to pour some coffee into the cup; can you set up the cup on the table so I can do this properly?” (requires reorienting the cup so that it is placed ‘right-side-up’ on the table). The optimal solution to this task requires a sequence of2 skills.
  • $\bullet$
    Mug1: “Setup the mug so it’s upright, then put whatever object that fits inside of it”. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    Mug2: “Place cutlery inside the mug and then place the mug itself on the table near the condiment” (the mug’s opening is obstructed by a large orange, which must be moved out of the way). The optimal solution to this task requires a sequence of 8 skills.
  • $\bullet$
    Mug3: Same as CutleryInMug2, except in the initial state the mug is not obstructed by an orange, but rather has a ball stuck inside it. This must be removed by ‘pouring’ it out of the cup itself). The optimal solution to this task requires a sequence of 8 skills.
  • $\bullet$
    SoupPour: “Serve the fruits on the white mat (make sure the peach is to the right of the apple” and pour soup into the red container”. The white mat is originally obstructed by the soup can as well as a spam can, which must be moved out of the way to successfully place the fruits. The optimal solution requires a sequence of 10 skills.
Approaches. We list the various approaches we compare to OWL-TAMP across various tasks.
  • $\bullet$
    CaP: A version of Code as Policies [35] adapted to our domain, and using a VLM instead of an LLM as in the original work. This approach attempts to query a VLM to directly specify a function that outputs a full plan given an initial state. We then execute each action directly without any sampling. We provide the VLM with the same few-shot examples, as well as all the same helper functions that our approach uses.
  • $\bullet$
    No VLM: Inspired by Liu et al. [57] and Xie et al. [13], we simply prompt an LLM to translate the given natural language goal to a symbolic goal expression, and call a TAMP system to achieve this goal.
  • $\bullet$
    CaP-sample: A version of CaP inspired by the method of Curtis et al. [7] where the VLM generates a function that generates a variety of different plans (where each plan differs in the continuous parameters provided to the skill). We sample plans from this function and return the first full plan that does not violate any robot constraints.
  • $\bullet$
    No sample: A version of OWL-TAMP where we only allow one continuous sample per skill in the plan.
  • $\bullet$
    No cont.: A version of OWL-TAMP where we ablate the continuous constraint generation module (i.e., all continuous constraints come directly from the underlying TAMP system) but enable discrete constraint generation.
  • $\bullet$
    No disc.: A version of OWL-TAMP where we ablate the discrete constraint generation module but enable continuous constraint generation.
  • $\bullet$
    No backtrack: A version of OWL-TAMP where we ablate the underlying TAMP system’s ability to backtrack over different task skeletons (i.e., it is confined to attempting to refine the first plan skeleton it finds).
Experimental Setup. We use GPT-4o [1] as our VLM for all approaches. We report results over 10 independent random seeds for each task; we keep the natural language goal the same, and randomize the poses of a predetermined number of objects. For ‘CaP-sample’, we provide a sampling budget of $2500$ samples from its generated LMP for each task. For all other methods (except ‘CaP’ and ‘No sample’, where the sampling budget is set to $1$), we provide a sampling budget of $500$ per action in the plan. For methods that perform backtracking (i.e., all variants of our approach except ‘No back.’), we provide a backtracking budget of $5$ (i.e., we allow the method to attempt $5$ distinct plan skeletons). Our main metric of interest is each method’s success rate across tasks. We consider that a particular method has succeeded at a task if the robot does not violate any environmental constraints, and achieved the natural-language goal as judged by manually-written task-specific success detector functions. We also measure and report the ‘soundness rate’, which indicates how often an approach produces false positives (i.e., it believes it has accomplished a task, but actually has not). We report and analyze additional metrics on this domain, such as the total the wall-clock time, number of plan skeletons, and number of calls to a pretrained model to find a solution, in Appendix A-C. For all methods in the RAVENS-YCB Manipulation domain, we use a common TAMP system based on the search-then-sample TAMP system widely used in previous work [8, 12, 16, 17].
Results and Analysis. Table I shows our main success rate results for the RAVENS-YCB Manipulation domain. OWL-TAMP achieves the highest success rate of all methods in 9 out of the 10 presented tasks. The ‘No VLM’ baseline performs well in the simplest 3 tasks, where the natural language goal can be exactly translated into a goal expression in terms of predicates available to the underlying TAMP system, but fails almost completely in the remaining 7 tasks. Both ‘CaP’ and ‘No sample’ fail across all tasks. Upon inspection, we found that the discrete plan skeletons produced by these methods were generally correct, but that they were unable to select continuous parameters to satisfy the various continuous constraints within the domain. ‘CaP-sample’ performs better than ‘CaP’ due to its ability to sample different continuous values until it finds a satisfying set, but fails in tasks where the goal expresses non-trivial constraints because it does not explicitly attempt to generate and solve for these constraints. The ‘no cont.’ and ‘no disc.’ ablations of OWL-TAMP succeed in the first 3 tasks that do not require any discrete or continuous constraints beyond those embedded in the TAMP system, but struggle beyond these. Finally, the ‘no back.’ ablation of OWL-TAMP succeeds in most tasks except the final 4, which tend to require moving obstacles out of the way in order to satisfy constraints. We find that the VLM is able to generate appropriate discrete constraints to enable this in short-horizon tasks without much clutter (e.g. Berry2), but requires the assistance of a TAMP system equipped with backtracking to handle longer-horizon tasks featuring more objects.
[…]
Figure 4 shows our main soundness rate results. These results demonstrate that OWL-TAMP generally does not produce false positives; in fact, there was only a single false positive across all tasks. This indicates that our approach is generally able to accurately capture the constraints within the various natural language goal expressions. By contrast several baselines — most notably ‘No VLM’ and ‘No cont.’ — have much higher satisfying plan rates than success rates (i.e., they falsely declare they have solved tasks). This is because they translate the natural language goal into problems that they solve, but that do not actually correspond to achieving the true underlying task goal.
Our approach failed at least once in ‘BerryCook’, ‘Mug3’, and ‘SoupPour’. Upon inspection, we found that all the failures in ‘BerryCook’ were due to incorrect continuous constraints. The VLM incorrectly constrained the strawberry to be simultaneously inside the bowl and pan at the end of the task in order to achieve the goal. In the ‘Mug3’ task, we found that our approach generated correct constraints, but was unable to find a satisfying set of samples with the provided budget. In the ‘SoupPour’ task, we found that the VLM generated incorrect discrete constraints: the TAMP system was not able to find a plan that included the generated partial plan as a sub-sequence.
Real-World System Deployment To test our approach on real-world hardware, We deployed OWL-TAMP on a custom dual-arm manipulator with two Kinova Gen3 arms and a pan-tilt head camera to perform a variety of natural language manipulation tasks. We demonstrated $19$ different tasks covering a wide range of generalized pick-and-place problems (i.e., each task was solvable via a specific pick-and-place sequence with appropriate continuous parameters) within this domain, $2$ of which are depicted in Figure 3. The goal descriptions for these tasks are listed below.
  • $\bullet$
    “Put the orange and apple on the plate.”
  • $\bullet$
    “Place the strawberry and lime each in the bin that matches their color.”
  • $\bullet$
    “Stack the blocks into a tower by increasing hue.”
  • $\bullet$
    “Put the apple left of the plate and the orange on the table surface behind of the plate.”
  • $\bullet$
    “Put the orange on the far right of the table and the apple on the far left”.
  • $\bullet$
    “Put the orange where the apple is initially”.
  • $\bullet$
    “Clean the plate” (a sponge is among several other objects present on a tabletop, and the robot must put the sponge atop the plate)
  • $\bullet$
    “Throw away anything not vegan in the purple bin” (objects on the table include a milk carton, apple, spam can, and water bottle).
  • $\bullet$
    “Put the green block between the blue and red ones”
  • $\bullet$
    “Put the blue block onto the plate” (the plate is packed with distractors and the robot must make a tightly-constrained placement).
  • $\bullet$
    “Setup the cutlery for someone to eat a meal from the plate. All the cutlery should be close to and lined-up with the plate, and should be oriented so each is straight and facing forwards, though you should pick which side of the plate each of the items are on” (there are two pieces of fruit, and two similarly-colored blocks that must be disambiguated).
  • $\bullet$
    “Fit one of the fruit in the cup” (only one of 4 available fruits is small enough to fit in the provided mug)
  • $\bullet$
    “Put the brownie ingredients in front of the pan” (only 2 of the available items are related to brownies)
  • $\bullet$
    “Place the cutlery in the utensil holder. All the cutlery should be oriented straight and facing forward”
  • $\bullet$
    “Fry two eggs at the front of the pan”
  • $\bullet$
    “Fry the spam on the pan and serve it on the plate”
  • $\bullet$
    “Weigh the shortest object and put it in the bin”
  • $\bullet$
    “Put the banana near the other fruit”
  • $\bullet$
    “Place the red block so that it’s aligned with the other two blocks”
Tasks were implemented by placing a set of objects in front of the robot and providing a natural language command. The robot leveraged a vision pipeline described in Appendix A-D to estimate the geometric scene, performed constraint generation and planning using this simulation, and then executed generated plans on the real hardware open-loop. We used the PDDLStream [9] system as the underlying TAMP system within OWL-TAMP. Execution videos of these tasks are included in our supplementary materials.
Overall, OWL-TAMP was successfully able to generate the necessary constraints to enable generally correct behavior in all these tasks, given just direct camera input and natural language goal descriptions. Importantly, we used the same system with no variation — even in prompting or the few-shot examples provided  [^4] We provided the same few-shot examples as used in our simulated tasks. — to accomplish this variety of tasks. Nearly all of the tasks required specific choices for continuous parameters (e.g. side grasps instead of purely top-down, constrained placements due to collisions and reachability), and several required the TAMP system to augment the VLM-proposed partial plan to move one or more objects out of the way. Several tasks included specific instructions that the TAMP system could not possibly have achieved given its internal predicates and operators and thus relied on the VLM-inferred constraints. Thus, our system’s ability to solve these tasks required close interplay between both the TAMP system and the VLM central to OWL-TAMP.
VI Limitations and Future Work
Our system relies entirely on the VLM to generate constraints that are both syntactically and semantically correct. There is currently no provision for recovering from errors — either syntactic or semantic — in the generated constraints. It is possible for the VLM to generate constraints — especially continuous constraints — for a task that are either not satisfiable at all, or that are very challenging to satisfy, which would cause the TAMP system to fail to find a solution. Finally, we assume a set of primitive parameterized robot skills (e.g. moving, rigid grasping, placement), which plans ultimately consist of. We also assume a library of primitive helper functions that the VLM leverages to construct continuous constraints, and we observed that generated constraints are usually simple compositions of helper functions.
Avenues for future work include enabling a feedback loop between the TAMP system and the VLM so that the VLM can modify or augment constraints if the TAMP system declares a problem infeasible, enabling the VLM to generate continuous parameter samplers to help the TAMP system satisfy generated continuous constraints, finetuning VLMs [58] on a representative distribution of robot manipulation tasks and constraints to improve their accuracy and generality at constraint generation, and investigating extensions of our framework to handling tasks with more complex (e.g. temporal) constraints.
VII Conclusion
We presented OWL-TAMP, a system that uses VLMs to infer discrete and continuous constraints that enable a TAMP system to interpret and accomplish open-world language instructions. We demonstrated that our approach supports multiple underlying TAMP systems, and is able to solve a wide range of complex, long-horizon manipulations tasks across different robot embodiments in simulation and the real world.
Acknowledgements
Work for this project was partially conducted at NVIDIA Research. We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grants N00014-22-1-2740 and N00014-24-1-2603; from ARO grant W911NF-23-1-0034; from the MIT Quest for Intelligence; and from the Boston Dynamics Artificial Intelligence Institute.
Appendix A Appendix
A-A Additional RAVENS-YCB Manipulation Task Details
Figure 5: Ravens YCB-Manipulation tasks visualization. From top left to bottom right: visualizations of the goal string and an example initial state for the ‘Berry1’, ‘Citrus’, ‘Berry2’, ‘BerryCook’, ‘FruitSort’, ‘Coffee’, ‘Mug1’, ‘Mug2’, ‘Mug3’, and ‘SoupPour’ tasks.
[…]
  • $\bullet$
    Berry1: Goal description: “put the strawberry onto the light-grey region at the center of the table”. Objects: strawberry, light_grey_region, table_surface. The position of the light_grey_region is held constant at the center of the table. The optimal solution to this task requires a sequence of 2 skills.
  • $\bullet$
    Citrus: Goal description: “pack the citrus fruit onto the plate”. Objects: strawberry, apple, pear, lemon, orange, plum, table_surface. The plate object is always set to be at the center of the table in the initial state. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    Berry2: Goal description: “put the strawberry onto the light-grey region at the center of the table”. Objects: strawberry, light_grey_region, potted_meat_can, table_surface. The only object whose position is randomized in the initial state is the strawberry: the potted_meat_can is always set to totally obstruct the surface of the light_grey_region in the center of the table: it must be moved out of the way to make room to place the strawberry and successfully accomplish the goal description. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    BerryCook: Goal description: “Cook the strawberry by putting it in the pan, then finally simply place it in the bowl. The strawberry should only be in the bowl at the end!”. Objects: strawberry, skillet, bowl, table_surface. The only object whose position is randomized in the initial state is the strawberry. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    FruitSort: Goal description: “Put all the fruit to the left of the line bisecting the table”. Objects: pear, sugar_box, hammer, tomato_soup_can, strawberry, apple, red_line, table_surface. The position of the red_line object is kept constant at the center of the table. The optimal solution to this task requires a sequence of 6 skills.
  • $\bullet$
    Coffee: Goal description: “I want to pour some coffee into the cup; can you set up the cup on the table so I can do this properly?” (requires reorienting the cup so that it is placed ‘right-side-up’ on the table). Objects: mug, table_surface. The optimal solution to this task requires a sequence of 2 skills.
  • $\bullet$
    Mug1: Goal description: “Setup the mug so it’s upright, then put whatever object that fits inside of it”. Objects: fork, power_drill, potted_meat_can, mug, table_surface. The initial position of the mug is fixed. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    Mug2: Goal description: “Place cutlery inside the mug and then place the mug itself on the table near the condiment” (the mug’s opening is obstructed by a large orange, which must be moved out of the way). Objects: fork, knife, sponge, strawberry, orange, mustard_bottle, mug, table_surface. The initial position of the mug and orange are fixed such that the orange is always obstructing the mouth of the mug. The optimal solution to this task requires a sequence of 8 skills.
  • $\bullet$
    Mug3: Goal description: “place cutlery into the mug and have the mug be near the condiment”. This is the same as CutleryInMug2, except in the initial state the mug is not obstructed by an orange, but rather has a ball stuck inside it. This must be removed by ‘pouring’ it out of the cup itself). Objects: fork, golf_ball, mustard_bottle, mug, table_surface. The initial position of the mug and golf_ball are fixed such that the white is always inside the mug. The optimal solution to this task requires a sequence of 8 skills.
  • $\bullet$
    SoupPour: Goal description: “Serve the fruits on the white mat (make sure the peach is to the right of the apple” and pour soup into the red container”. The white mat is originally obstructed by the soup can as well as a spam can, which must be moved out of the way to successfully place the fruits. Objects: white_mat, tomato_soup_can, potted_meat_can, bowl, apple, peach, table_surface. The only object positions that are randomized in the initial state are that of the apple and the peach. The optimal solution to this task requires a sequence of 10 skills.
A-B Additional Method Implementation Details
Here, we first provide a detailed explanation of how we implement the TAMP system used in our simulation experiments (Appendix  A). Subsequent sections (Appendix A-E, Appendix A-F, and Appendix A-G) provide more specific details around prompting. For details beyond those presented here, we refer the reader to our open-source code release, which we will make available after acceptance.
A-B1 TAMP System Details
We implement a relatively simple variant of the search-then-sample TAMP approach used in several recent works [8, 12, 16, 17] and use this as the underlying TAMP system for all methods that require one (i.e., all methods except ‘CaP’ and ‘CaP-sample’). We implement versions of the same predicates and operators described in Section III (including a version of transport that we call Pour), though we omit the move operator (movement is performed automatically as part of attach and detach) choose to name them differently (as illustrated by our prompts in Appendix A-F below). We manually define the initial state of all tasks in terms of literals involving these predicates; the robot always begins each task at the same initial configuration, and with the HandEmpty() predicate set to True. Additionally, we associate each operator with a natural language description to make it easier for the VLM to perform discrete constraint generation (e.g., for the detach operator, the description is something like: “places object o stably atop a surface; you can specify a description of how this placement should be performed by filling in the description parameter accordingly”). Each operator is linked to one particular low-level skill included with the environment.
We define an associated sampler for each of these operators. Each sampler takes in the current state of the task, as well as the operator’s discrete arguments, and defines a distribution over the operator’s continuous parameters. These samplers are setup to define broad distributions (e.g. the sampler for the detach operator simply tries to find a pose somewhere broadly above the surface to be detached onto, but roughly within the confines of the x and y boundaries of the surface’s axis-aligned bounding box), though we do modify these slightly depending on the task (e.g. the sampler for the ‘Coffee’ and ‘Mug’ tasks defines particular distributions on the orientations the mug and/or cutlery should be Placed).
Given these components, we can adopt the following hierarchical planning strategy to achieve a symbolic predicate goal $G$ from initial state $s_{0}$ 555see [12] or [16] for a fuller presentation of the search-then-sample TAMP strategy we employ.:
  1. 1.
    Compute a new task plan (sequence of ground operators) that achieves $G$ from $s_{0}$
  2. 2.
    For $b$ within the backtracking budget:
    1. (a)
      For each ground operator in this plan:
      1. i.
        If the preconditions do not hold, break.
      2. ii.
        For $i$ within the sampling budget:
        1. A.
          Call the associated sampler to get continuous parameters.
        2. B.
          Use these to instantiate the skill associated with the operator.
        3. C.
          Execute the skill within the planning model (i.e., the simulator)
        4. D.
          Check that the resulting state satisfies all the predicate effects.
Our task planning is performed initially via a simple $A^{*}$ search. During backtracking (i.e., when the sampling budget is exhausted for the first time and a new task plan is required), we employ a set of manually-engineered strategies to modify the task plan based on the most-recent failed operator (e.g. if the most recent-failed operator is a detach that was attempting to place an object atop a particular surface, and there are other objects atop that surface already, we randomly append a attach detach sequence to move one of those objects to a different part of the table).
A-B2 OWL-TAMP Constraint Generation Example Walkthrough
Below walk through a concrete example of generating both discrete and continuous constraints for the ‘Mug1’ task.
First, we prompt the model to generate discrete constraints given the task goal. It generates the following partial plan, along with language descriptions filled in for each action:
1(’Pick(mug)’, ’grasp the mug securely to lift it from the table.’)
2(’Place_Ontop(mug, table)’, ’place the mug upright on the table to ensure it is stable.’)
3(’Pick(fork)’, ’grasp the fork securely to lift it from the table.’)
4(’Place_Inside(fork, mug)’, ’carefully place the fork inside the upright mug.’)
5(’achieve_goal(mug, fork)’, ’the goal is achieved when the mug is upright and the fork is inside it.’)
Next, we prompt the model to provide continuous constraints for the achieve_goal operator at the end of the plan. This corresponds to providing continuous constraints that correspond to the task’s goal description.
1Goal Check Functions:
2def goal_check0() -> bool:
3 upright_mug = abs(mug.pose.roll) < 0.1 and abs(mug.pose.pitch) < 0.1 and abs(mug.pose.yaw) < 0.1
4 return upright_mug
5
6
7def goal_check1() -> bool:
8 inside_mug_bounds = modify_pose_bounds_to_be_ inside_object(init_state, env, init_bounds, mug.category)
9 return position_within_bounds(fork.pose, inside_mug_bounds)
Next, we prompt the model to generate continuous constraints for each operator given these generated constraint functions as input. We start from the beginning of the plan. The Pick(mug) operator (which corresponds to attach) does not have any effects or constraint predicates that depend on the VLM, so we skip this operator. The Place_Ontop(mug, table) operator ((which corresponds to detach) does feature a VLMPose constraint. So we prompt the VLM to generate continuous constraints that implement it based on the language description as well as the goal generated constraints.
1Goal Check Functions:
2def goal_check2() -> bool:
3 upright_mug = abs(mug.pose.roll) < 0.1 and abs(mug.pose.pitch) < 0.1 and abs(mug.pose.yaw) < 0.1
4 on_table_bounds = modify_pose_bounds_to_ be_ontop _of_object(init_state, env, init_bounds, mug.category, ’table’)
5 return upright_mug and position_within_bounds(mug.pose, on_table_bounds)
The Place_Inside(fork, mug) operator (which also corresponds to detach) does not have any effects or constraint predicates that depend on the VLM, so we skip this operator as well. Finally, the Place_Inside(fork, mug) does have a VLMPose constraint, so we once again prompt the VLM, and obtain the following continuous constraints for this operator:
1Goal Check Functions:
2def goal_check2() -> bool:
3 inside_mug_bounds = modify_pose_bounds_ to_be_inside_object(init_state, env, init_bounds, mug.category)
4 return position_within_bounds(fork.pose, inside_mug_bounds)
Note that in this particular case, the VLM only checks that the fork is inside the mug, which is a constraint that is built into the TAMP system itself (via the Inside predicate that is an effect of the Place_Inside operator).
Now, we have a partial plan along with continuous constraints for particular actions. We call our underlying TAMP system to attempt to satisfy these, as well as its own internal constraints.
A-B3 Code as Policies Implementation Details
Following recent work [7], we adapt Code as Policies to our RAVENS YCB-Manipulation domain by having it output a Language Model Program (LMP) function that — given an initial state — directly outputs a full plan with skills and continuous parameters. We provide it with all the same helper functions used for continuous parameter generation that our method has access to. We also provide it with few-shot example solutions to the same 3 example problems as we provide in the continuous constraint generation prompt of our approach. Additionally, we provide it with one sampler per skill that it can use to generate continuous parameters for each respective skill. We provide the full detailed prompt we use for this method below in Appendix A-G.
Tasks
Method Berry1 Citrus Berry2 BerryCook FruitSort Coffee Mug1 Mug2 Mug3 SoupPour
CaP-sample 83.30 $\pm$110.55 2185.30 $\pm$546.48 2353.50 $\pm$363.06 2606.00 $\pm$25.00 0.00 $\pm$0.00 255.60 $\pm$161.55 2519.90 $\pm$3.45 500.90 $\pm$755.41 2557.10 $\pm$9.09 1770.70 $\pm$874.10
No VLM 34.70 $\pm$19.45 247.70 $\pm$81.38 619.40 $\pm$28.15 15.90 $\pm$13.46 0.00 $\pm$0.00 0.00 $\pm$0.00 856.40 $\pm$661.17 819.50 $\pm$374.59 320.20 $\pm$186.27 113.50 $\pm$57.43
No disc. 42.10 $\pm$16.74 303.70 $\pm$85.50 729.30 $\pm$126.84 47.00 $\pm$28.48 0.00 $\pm$0.00 0.00 $\pm$0.00 1636.90 $\pm$349.74 1293.00 $\pm$208.14 915.30 $\pm$244.54 208.70 $\pm$139.25
No cont. 35.10 $\pm$19.37 349.40 $\pm$244.23 193.50 $\pm$115.46 41.00 $\pm$25.54 308.30 $\pm$167.53 132.40 $\pm$78.32 376.80 $\pm$299.85 2500.00 $\pm$0.00 1478.70 $\pm$512.81 1170.60 $\pm$362.64
No back. 37.70 $\pm$16.59 229.10 $\pm$105.03 181.20 $\pm$109.62 36.40 $\pm$17.54 348.20 $\pm$86.87 167.60 $\pm$77.11 348.30 $\pm$145.08 369.50 $\pm$150.42 500.00 $\pm$0.00 450.00 $\pm$113.11
OWL-TAMP 37.70 $\pm$16.59 250.30 $\pm$140.93 181.20 $\pm$109.62 101.90 $\pm$67.60 372.40 $\pm$114.46 167.60 $\pm$77.11 696.30 $\pm$508.87 710.00 $\pm$461.70 1296.30 $\pm$607.64 1335.10 $\pm$409.67
Manual 42.10 $\pm$16.74 303.70 $\pm$85.50 120.60 $\pm$50.45 43.30 $\pm$12.77 271.70 $\pm$99.02 197.80 $\pm$125.70 365.10 $\pm$312.24 657.70 $\pm$666.77 264.40 $\pm$202.57 832.80 $\pm$785.73
TABLE II: Number of samples on all tasks. We report the average number of continuous parameter sampling attempts for each task across $10$ random seeds; the number of samples within the 95% confidence interval is reported after the $\pm$ sign for every entry. We also include a row (‘Manual’) reporting results on a variant of OWL-TAMP where we manually specify ground-truth discrete and continuous constraints to serve as an approximate lower bound on the number of samples our approach could achieve.
Tasks
Method Berry1 Citrus Berry2 BerryCook FruitSort Coffee Mug1 Mug2 Mug3 SoupPour
CaP 10.03 $\pm$0.32 (47.38%) 21.27 $\pm$1.10 (25.02%) 12.63 $\pm$1.13 (42.76%) 13.77 $\pm$0.95 (45.88%) 19.50 $\pm$0.67 (26.14%) 9.66 $\pm$0.80 (47.61%) 19.77 $\pm$2.84 (46.16%) 23.25 $\pm$1.65 (30.42%) 16.37 $\pm$1.49 (38.14%) 19.33 $\pm$1.17 (42.88%)
CaP-sample 22.40 $\pm$12.70 (21.22%) 248.49 $\pm$54.38 (2.14%) 245.31 $\pm$35.87 (2.20%) 262.43 $\pm$14.12 (2.41%) 20.88 $\pm$2.05 (24.41%) 40.39 $\pm$16.21 (11.39%) 281.74 $\pm$2.99 (3.24%) 75.61 $\pm$78.73 (9.35%) 274.50 $\pm$11.97 (2.27%) 187.90 $\pm$83.47 (4.41%)
No VLM 11.90 $\pm$1.35 (8.85%) 41.59 $\pm$4.85 (4.53%) 47.66 $\pm$3.20 (1.60%) 15.96 $\pm$1.23 (29.77%) 16.56 $\pm$0.14 (24.60%) 4.92 $\pm$0.04 (10.25%) 59.43 $\pm$31.91 (4.11%) 78.72 $\pm$18.85 (6.36%) 36.03 $\pm$11.85 (8.04%) 21.28 $\pm$2.54 (20.83%)
No sample 44.00 $\pm$4.03 (85.97%) 75.17 $\pm$5.26 (72.29%) 59.47 $\pm$4.69 (83.77%) 54.93 $\pm$4.68 (82.13%) 81.19 $\pm$10.40 (74.17%) 39.95 $\pm$4.28 (85.49%) 55.07 $\pm$3.43 (73.59%) 80.00 $\pm$9.63 (69.38%) 46.22 $\pm$2.57 (70.58%) 91.46 $\pm$29.02 (82.12%)
No disc. 44.92 $\pm$4.07 (70.03%) 91.62 $\pm$6.22 (50.92%) 80.87 $\pm$6.83 (33.37%) 44.50 $\pm$4.38 (69.50%) 36.10 $\pm$2.01 (59.87%) 18.92 $\pm$2.96 (72.29%) 118.42 $\pm$16.71 (20.71%) 120.94 $\pm$9.99 (29.63%) 81.85 $\pm$10.58 (26.85%) 96.95 $\pm$6.37 (72.54%)
No cont. 18.15 $\pm$1.84 (40.92%) 54.53 $\pm$13.09 (16.21%) 36.91 $\pm$5.61 (29.48%) 24.30 $\pm$1.91 (29.40%) 57.03 $\pm$9.88 (15.68%) 25.73 $\pm$4.13 (29.93%) 52.26 $\pm$16.08 (14.61%) 136.88 $\pm$5.72 (5.72%) 102.53 $\pm$22.20 (6.04%) 91.00 $\pm$20.21 (9.42%)
No back. 48.69 $\pm$3.40 (71.92%) 106.36 $\pm$24.85 (60.35%) 101.84 $\pm$34.28 (72.59%) 67.94 $\pm$15.09 (75.11%) 149.53 $\pm$39.78 (66.52%) 57.68 $\pm$8.20 (65.30%) 87.31 $\pm$21.78 (58.23%) 106.82 $\pm$6.13 (54.49%) 69.59 $\pm$3.02 (52.24%) 111.50 $\pm$32.13 (67.36%)
OWL-TAMP 46.92 $\pm$3.37 (74.64%) 108.50 $\pm$24.61 (59.15%) 101.88 $\pm$34.47 (72.56%) 71.38 $\pm$15.31 (71.49%) 154.32 $\pm$43.43 (64.46%) 57.96 $\pm$7.69 (64.98%) 113.18 $\pm$28.50 (44.92%) 140.51 $\pm$23.06 (41.43%) 118.38 $\pm$23.74 (30.71%) 173.81 $\pm$49.00 (43.21%)
Manual 13.53 $\pm$2.36 (0.00%) 44.42 $\pm$5.34 (0.00%) 24.84 $\pm$2.78 (0.00%) 19.78 $\pm$1.34 (0.00%) 49.72 $\pm$3.59 (0.00%) 22.15 $\pm$6.36 (0.00%) 45.18 $\pm$16.74 (0.00%) 79.91 $\pm$33.04 (0.00%) 37.64 $\pm$10.40 (0.00%) 81.38 $\pm$41.85 (0.00%)
TABLE III: Wall clock time on all tasks. We report the average wall clock time in seconds each approach took to solve each task across $10$ random seeds; the time within the 95% confidence interval is reported after the $\pm$ sign, and the average percentage of that time that was spent querying a foundation model is reported within parentheses ’()’ for every entry. We also include a row (‘Manual’) reporting results on a variant of OWL-TAMP where we manually specify ground-truth discrete and continuous constraints (and thus no need to query a foundation model for these) to serve as an approximate lower bound on the time our approach could achieve.
A-C Additional Experimental Results
In this section, we explore experimental results related to how much computation and time our various methods took on each of the simulated tasks in the RAVENS YCB-Manipulation domain from Section V. We also present a fine-grained table of soundness rate results used to construct Figure 4.
Table II shows the number of samples required by the various methods to solve each of the RAVENS-YCB Manipulation tasks. Note importantly that we do not adjust for success rates (depicted in Table I) here: many approaches that achieve a low number of samples (e.g. CaP-sample or No VLM on FruitSort) here actually do so only because they in fact fail to translate that particular task into a form that they can attempt to solve via sampling. OWL-TAMP generally does not require significantly more samples than the ‘Manual’ baseline (which serves as an approximate lower bound on the samples). In two of the tasks where it does require substantially more samples than ‘Manual’ (Mug3, SoupPour), OWL-TAMP provides discrete constraints that yield an initial plan skeleton that is incorrect, necessitating backtracking within the TAMP system, which significantly increases the number of samples.
Table III shows the wall clock time required by the various methods to solve each of the RAVENS-YCB Manipulation tasks. Note once again that we do not adjust for success rates here. In general, we see that our approach takes significantly more wall clock time than the ‘Manual’ oracle, but we also see that a substantial portion of the time taken (around 50% on average) is due to querying a foundation model. We also see that the percentage of time spent querying increases roughly with the task complexity (left to right), which is what we would expect, since in the more complex tasks with more constraints, it is more challenging to find a satisfying sample. These findings indicate that the overall runtime of our method could be significantly reduced by reducing the time taken for foundation model querying.
Tasks
Method BerryCook FruitSort Coffee Mug1 Mug2 Mug3 SoupPour
CaP-sample 100% 100% 0% 100% 100% 100% 100%
No VLM 0% 0% 0% 20% 0% 20% 100%
No disc. 40% 100% 90% 100% 100% 100% 100%
No cont. 100% 10% 60% 70% 100% 20% 20%
OWL-TAMP 100% 100% 100% 100% 100% 100% 90%
[…]
A-D Real Robot System Implementation
For the real-world demos, we adapt the strategy of Curtis et al. [59], which deploys TAMP without a priori object models by estimating collision, grasp, and placement affordances online. We deploy OWL-TAMP in a replanning policy [60], where at each state, the robot observes the world with its head camera, segments and estimates the object geometries, plans a course of action, and executes its plan. We used Grounding DINO [61] for object detection, segmentation, and association, RVT [62] to infer grasp affordances, TRAC-IK [63] for inverse kinematics, and cuRobo [64] for motion planning.
A-E Helper Functions for Continuous Constraint Generation.
We provide all methods (Section V) access to the following helper functions to be used towards constraining the continuous parameters used to instantiate skills. We provide the name, parameters and docstring of each function exactly as below:
1def get_aabb_bounds
2 """Given the state of a particular env, and an object_name that appears in
3 this state, return tuples corresponding to the bounds of the axis-aligned
4 bounding box of object_name in this state in the world frame.
5
6 In particular, return the lower and upper bounds on the axis-aligned
7 x, y, z values.
8 """
9
10def get_obj_center
11 """Given the state of a particular env, return the pose of the object with
12 object_name.
13
14 The pose is a tuple of dim 6 corresponding to (x, y, z, roll, pitch,
15 yaw).
16 """
17
18def modify_pose_bounds_to_be_behind_object
19 """Given a tuple of initial bounds (init_bounds), return a modified set of
20 bounds such that sampling randomly from the output bounds will ensure that
21 a pose will be selected that is behind (on the table plane) the object with
22 name ‘object_name‘’s. For instance:
23 modify_pose_bounds_to_be_behind _object(init_state, env, init_bounds,
24 ’hammer’) will modify init_bounds such that they only contain poses that
25 are behind the ’hammer’ object on the table surface ahead the robot.
26
27 Note that this does not constrain
28 the pose’s horizontal position (it may be anywhere on the table - in the
29 left or right half - such that it’s behind object_name).
30 """
31
32def modify_pose_bounds_to_be_in_front_of_object
33 """Given a tuple of initial bounds (init_bounds), return a modified set of
34 bounds such that sampling randomly from the output bounds will ensure that
35 a pose will be selected that is to the in front of (on the table plane) the
36 object with name ‘object_name‘’s. For instance:
37 modify_pose_bounds_to_be_in_front_of_ object(init_state, env, init_bounds,
38 ’hammer’) will modify init_bounds such that they only contain poses that
39 are in front of the ’hammer’ object on the table surface ahead the robot.
40
41 Note that this does not constrain
42 the pose’s horizontal position (it may be anywhere on the table - in the
43 left or right half - such that it’s in front of object_name).
44 """
45
46def modify_pose_bounds_to_be_left_of_object
47 """Given a tuple of initial bounds (init_bounds), return a modified set of
48 bounds such that sampling randomly from the output bounds will ensure that
49 a pose will be selected that is to the left of (on the table plane) the
50 object with name ‘object_name‘’s. For instance:
51 modify_pose_bounds_to_be_left _of_object(init_state, env, init_bounds,
52 ’hammer’) will modify init_bounds such that they only contain poses that
53 are to the left of the ’hammer’ object on the table surface ahead the
54 robot.
55
56 Note that this does not constrain
57 the pose’s vertical position (it may be anywhere on the table - in the
58 upper or lower half - such that it’s to the left of object_name).
59 """
60
61def modify_pose_bounds_to_be_right_of_object
62 """Given a tuple of initial bounds (init_bounds), return a modified set of
63 bounds such that sampling randomly from the output bounds will ensure that
64 a pose will be selected that is to the right of (on the table plane) the
65 object with name ‘object_name‘’s. For instance:
66 modify_pose_bounds_to_be_right_ of_object(init_state, env, init_bounds,
67 ’hammer’) will modify init_bounds such that they only contain poses that
68 are to the right of the ’hammer’ object on the table surface ahead the
69 robot.
70
71 Note that this does not constrain
72 the pose’s vertical position (it may be anywhere on the table - in the
73 upper or lower half - such that it’s to the right of object_name).
74 """
75
76def modify_pose_bounds_to_be_above_object
77 """Given a tuple of initial bounds (init_bounds), return a modified set of
78 bounds such that sampling randomly from the output bounds will ensure that
79 a pose will be selected that is above (on the table plane) the object with
80 name ‘object_name‘’s. For instance:
81 modify_pose_bounds_to_be_ above_object(init_state, env, init_bounds,
82 ’hammer’) will modify init_bounds such that they only contain poses that
83 are above the ’hammer’ object on the table surface ahead the robot.
84
85 Note that this does actually also constrain the pose’s horizontal
86 position and vertical positions so that it is directly above the
87 object in question. Note also that this function might particularly
88 useful to constraint pouring actions (because pouring must be done
89 from above); though you will also likely have to apply an additional
90 angular constraint (since this function doesn’t apply any angular
91 constraints on its own).
92 """
93
94def modify_pose_bounds_to_be_below_object
95 """Given a tuple of initial bounds (init_bounds), return a modified set of
96 bounds such that sampling randomly from the output bounds will ensure that
97 a pose will be selected that is below (on the table plane) the object with
98 name ‘object_name‘’s. For instance:
99 modify_pose_bounds_to_be below_object(init_state, env, init_bounds,
100 ’hammer’) will modify init_bounds such that they only contain poses that
101 are below the ’hammer’ object on the table surface ahead the robot.
102
103 Note that this does actually also constrain the pose’s horizontal
104 position and vertical positions so that it is directly below the
105 object in question. Note also that this function might particularly
106 useful to constraint pouring actions (because pouring must be done
107 from above into a container that’s below); though you will also
108 likely have to apply an additional angular constraint (since this
109 function doesn’t apply any angular constraints on its own).
110 """
111
112def modify_pose_bounds_to_be_near_object
113 """Given a tuple of initial bounds (init_bounds), return a modified set of
114 bounds such that sampling randomly from the output bounds will ensure that
115 a pose will be selected such that the distance of the pose from the object
116 with name ‘object_name‘ will be within closeness_thresh along the x, y, and
117 z axes respectively (note that the pose might have an L2) distance that’s
118 greater than that."""
119
120def modify_pose_bounds_to_be_ontop_of_object
121 """Assuming the init_bounds are on the pose (x, y, z, roll, pitch, yaw) of
122 an object with name obj1_name, modify these such that the pose must be
123 confined to be on top of the object with name obj2_name.
124
125 Specifically, restrict the bounds to be within x and y of
126 obj2_name’s bounding-box, but have its z-position touching the top
127 of the bounding box of obj2_name.
128
129 IMPORTANT: use this only when trying to place an object atop another
130 (e.g. atop a region, or a surface of another object). If you want to put
131 something inside a container, use the
132 modify_pose_bounds_to_be_ontop_of_object function instead.
133 """
134
135def modify_pose_bounds_to_be_inside_object
136 """Assuming the init_bounds are on the pose (x, y, z, roll, pitch, yaw) of
137 an object with name obj1_name, modify these such that the pose must be
138 confined to be inside the object with name obj2_name.
139
140 Specifically, restrict the bounds to be within x and y of
141 obj2_name’s bounding-box.
142
143 IMPORTANT: use this only when trying to place an object inside a container
144 (e.g. a cup, or vase, or 3D box). If you want to put something in a 2D
145 region, use the modify_pose_bounds_to_be_ontop_of_object function instead.
146 Also note that this function is generally not suitable to constrain
147 pouring; it should generally be used when constraining placement!
148 """
149
150def position_within_bounds
151 """Checks that the xyz position component of a 6-d pose is within specific
152 bounds."""
153
154def initialize_bounds_anywhere_on_object
155 """Given obj, get its aabb and initialize bounds such that sampling within
156 these bounds will yield a pose with a position atop obj and any arbitrary
157 rotation."""
158
159def sample_ravenpose_uniformly_within_bounds
160 """Given obj, get its aabb and initialize bounds such that sampling within
161 these bounds will yield a pose with a position atop obj and any arbitrary
162 rotation."""
163
164def modify_obj_pose
165 """Modifies the pose of obj to new_pose."""
A-F OWL-TAMP prompting details.
As described in Section IV, our approach consists of an initial discrete constraint generation phase followed by continuous constraint generation based on a discrete partial plan skeleton with language parameters filled in.
We use separate prompts for both the discrete and continuous constraint generation. For both discrete and continuous constraint generation, we provide an image of the initial state of the task as part of the prompt.
Our discrete constraint generation prompt provides a single few-shot example of expected output on a particular task, and then requests the VLM to output a plan in a similar format for the current task. It also leverages chain of thought prompting [65] to encourage the model to improve the output accuracy. The full prompt we use is shown below: variables within curly brackets ({}) are filled-in dynamically depending on the task.
1You are an expert-level robot task planning system whose job is to help a robot accomplish the following task: ’’{task_str}’’.
2
3Here is the initial predicate state (i.e., the set of all ground atoms that are true) of this task. Note that an image corresponding to the environment in this state
4is attached below:
5{initial_preds}
6
7Your job is to output a sequence of ground operators (i.e., a plan) that ideally achieve the goal from this initial state.
8Your plan need not be perfect, but it should capture the critical objects and actions necessary to accomplish this task (e.g.
9if the task requires 4 objects being in a specific location, then you should take care to make sure the plan contains
10an action to manipulate each of the 4 objects in turn).
11
12Here are the unground operators with their descriptions.
13{nsrts_description}
14
15Here are all the ground operators available to you; each operator you use in your plan must be one of these.
16{ground_operators}
17
18Along with each operator in your plan, you should also output a natural language description of what that operator should
19do. This description can be as detailed as you like, and should explain any details relevant to completing the particular
20ground operator successfully.
21
22As an example, consider the example task ’’serve the banana inside the blue thing’’. Here, the bowl happens to be blue, and
23the initial state is:
24OnTable(banana)
25OnTable(bowl)
26And the available ground operators are:
27pick(banana)
28pick(bowl)
29pick(table)
30place_ontop(banana, bowl)
31place_inside(banana, bowl)
32place_ontop(bowl, banana)
33place_inside(bowl, banana)
34place_ontop(banana, table)
35place_inside(banana, table)
36place_ontop(bowl, table)
37place_inside(bowl, table)
38place_ontop(table, bowl)
39place_inside(table, bowl)
40
41Given this, the output should be something like:
42"""
43In the initial state, there is a blue bowl on the table, and a banana atop the table. The banana is not in the bowl, and the task is to
44move the banana into the bowl.
45The main actions relevant to the task are ‘pick(banana)‘ and ‘place(banana, bowl)‘. The goal involves a relationship between the banana and the bowl only.
46All other objects can be ignored.
47Plan:
48pick(banana); make a stable grasp on the banana - try to make a top-down grasp for maximum likelihood of success
49place(banana, bowl); place the banana stably so that it rests in the bowl - the banana is too large to fit inside the bowl if it is placed flatly: it needs to be reoriented to be upright so that it can fit into the bowl
50achieve_goal(banana, bowl); the goal involves the banana being inside the bowl - this relationship is purely between the banana and bowl and doesn’t involve/require any other objects.
51"""
52Notice how the plan ends in an achieve_goal operator. Every plan you output should end with such an operator, and the object arguments
53to this operator (i.e., ‘(banana, bowl)‘ in this case) should be all the objects necessary to decide whether or not the goal has been achieved
54(i.e., do your best not to include extraneous objects that are irrelevant to deciding whether the task goal has been achieved or not).
55
56
57Please output your plan in the following format (do not include the angle brackets: those are just for illustrative purposes). Importantly, please do not list the plan with a numbered or bulleted list,
58simply output each ground operator on a new line with no marking in front of the line as indicated below.
59<description of the initial state and task in your own words>
60<description of which objects and actions are particularly relevant to solving the task>
61<description of any challenges or other important considerations/obstacles that might arise when solving the task>
62Plan:
63<ground_operator0>; <natural language description0>
64<ground_operator1>; <natural language description1>
65...
66<ground_operatorm>; <natural language descriptionm>
Note that we ask the model to output a achieve_goal operator at the end of the plan. This is used by the continuous constraint generation procedure that follows: we generate continuous constraints for this action (the natural language description of the achieve_goal operator is effectively the task’s goal description $g$) first — which effectively corresponds to generating constraints for the task goal — and then generate constraints for any previous operators in the plan by conditioning on these goal continuous constraints.
We prompt the VLM to generate goal continuous constraints by providing the helper functions listed above, available objects in the scene, as well as three few-shot examples (shown below) of outputs on three separate simple example problems:
1To give you an idea of what your output function should look like, here is an example function generated for the task "put the lemon on the spoon and the banana on the table", where "lemon", "spoon", "banana", and "table" are all objects in that task/scene.
2
3‘‘‘python
4def goal_check0() -> bool:
5 ontop_spoon_bounds = modify_pose_bounds_to_ be_ontop_of_object(init_state, env, init_bounds, lemon.category, spoon.category)
6 return position_within_bounds(lemon.pose, ontop_spoon_bounds)
7‘‘‘
8‘‘‘python
9def goal_check1() -> bool:
10 ontop_table_bounds = modify_pose_bounds_to_ be_ontop_of_object(init_state, env, init_bounds, banana.category, table.category)
11 return position_within_bounds(banana.pose, ontop_table_bounds)
12‘‘‘
13
14Here is another example set of functions generated for the task "serve the banana inside the blue thing after drying it by placing on the plate". The initial state for this example
15is shown in one of the attached images. Here, banana and bowl are both objects (the bowl happens to be blue, whereas the plate is red).
16The initial state in this example is:
17bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
18banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
19plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
20table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
21
22Importantly, notice how the goal_check function checks that the banana is ’upright’ in the bowl by checking its rotation is 90 degrees (approx. 1.57 radians)
23along the roll axis. This is necessary, because the banana only fits into the bowl in this orientation, as shown in another example rendered
24image attached below. Pay careful attention and think about any similar orientation constraints that might be necessary in new problems.
25
26‘‘‘python
27def goal_check0() -> bool:
28 in_bowl_bounds = modify_pose_bounds_to_ be_inside_object(init_state, env, init_bounds, bowl.category)
29 banana_in_bowl_bounds = position_within_bounds(banana.pose, in_bowl_bounds)
30 is_upright = 1.4 <= abs(banana.pose.roll) <= 1.65
31 return banana_in_bowl_bounds and is_upright
32‘‘‘
33Notice here that only one goal_check function was each defined, because satisfying the goal depends on all the continuous variables jointly.
34Notice also that the goal_check doesn’t test for anything to do with the plate, even though "drying" the banana in the plate was important to the task. This is because - in the final state - the banana should be in the bowl (it should have previously been placed in the cup), and the ‘goal_check‘ function only operates in the final state.
35
36Finally, here’s an example of constraints for a task "serve spam from its can into the cup". Here, the objects available are potted_meat_can and mug‘.
37‘‘‘python
38def goal_check0() -> bool:
39 above_mug_bounds = modify_pose_bounds_to_ be_above_object(init_state, env, init_bounds, mug.category)
40 above_mug = position_within_ bounds(potted_meat_can.pose, above_bowl_bounds)
41 pour_angle_sufficient = abs(potted_meat_can.pose.roll) > 1.2
42 return above_bowl and pour_angle_sufficient
43‘‘‘
44Notice once again that only one goal_check function was defined.
45Notice also that the function checks the roll of the potted_meat_can, because this is important to know that it has been sufficiently ’tipped-over’ such that its contents can fall from the bowl inside it into the cup.
46
47Carefully consider these examples to inform your own functions for the current problem.
48...
We use these same few-shot examples across all tasks without varying them. We prompt the model to output goal_check functions for the current goal and extract these. Then, for each of the previous operators in the plan with language descriptions as well as predicates that rely on the VLM for their implementation, we prompt the VLM for continuous constraints that should hold after that particular operator is executed.
A-G Code as Policies Full Prompt
Our full prompt for Code as Policies is shown below. Note that variables within curly brackets ({}) are filled-in dynamically depending on the task.
1You are operating in an environment that has access to the following classes: {env_structs}.
2
3You have access to the following variables. Note that TABLE_AABB.lower and TABLE_AABB.upper are tuples of length 3 that represent the lower and upper bounds on the x, y, and z positions that are on the table.
4{typed_variables}
5
6You also have access to the following objects in the environment.
7Each of these has type RavenObject‘. You are not able to use other objects:
8{object_names}
9
10Note that when you use/invoke these objects, do not use them as a string (i.e., do not provide quotes ’’). Instead
11use them directly (e.g. use banana, not ’banana’). Use the name of the object as a string only where explicitly required
12(e.g. by the GraspSampler).
13
14Here’s the poses of all the objects in the initial state of the scene (depicted in an attached image). Note that the axes of each object is shown as red (x-axis), green (y-axis), and blue (z-axis).
15Note that roll is rotation about the x-axis, pitch is rotation about the y-axis and yaw is rotation about the z-axis. Pay careful attention to the axes and the current orientations of objects in the initial state as and when you decide to write functions involving orientations.
16{init_state}
17
18
19You also have access to helper functions whose signatures and docstrings are shown below. Pay careful attention to the arguments and return values of each function.
20Helper function signatures:
21{helper_functions_and_docstrings}
22
23You have access to the following set of skills expressed as pddl predicates followed by descriptions.
24You have no other skills you can use, and you must exactly follow the number of inputs described below.
25The coordinate axes are x, y, z where x is distance from the robot base, y is left/right from the robot base, and z is the height off the table.
26
27Action("pick", [[o], g])
28Pick up object o at grasp g sampled from a grasp sampler. Note that you should use the object o directly and not use its name (e.g. banana and not banana’).
29
30Action("place_ontop", [[o, s], p])
31If holding an object o (e.g. ‘banana‘) place the object ontop surface s (e.g. ‘table‘) at pose p. Note that you should invoke s and o as objects directly and not use their names (e.g. banana and not banana’).
32
33Action("place_inside", [[o, c], p])
34If holding an object o (e.g. ‘banana‘) place the object inside container c (e.g. ‘bowl‘) at placement pose p. Note that you should invoke o and c as objects directly and not use their names (e.g. banana and not banana’).
35
36Action("pour", [[o, c], p])
37If holding an object ‘o‘ (e.g. ‘tomato_soup_can‘) pour from the object onto/inside of container ‘c‘ (e.g. ‘bowl‘) at a pouring pose dictated by p. Specifically, p is 4 numbers: <x, y, z> corresponding to the position to reach, and <theta> corresponding to the angle (in radians; -3.14 to 3.14) to tilt the hand to accomplish the pour. Note that you should invoke o and c as objects directly and not use their names (e.g. banana and not banana’).
38
39Finally, to help you generate continuous parameters required for these skills, you have access to the following samplers.
40
41@dataclass
42class GraspSampler(Sampler):
43 curr_state: RavenState
44 env: RavenYCBEnv
45 object_name: str
46
47 def sample(self, rng: np.random.Generator) -> List[float]:
48 """Simply return some value within the AABB of the object, and at any
49 orientation."""
50 aabb_lb, aabb_ub = get_aabb_bounds(
51 self.curr_state, self.env, self.object_name
52 )
53 lb = (aabb_lb[0], aabb_lb[1], aabb_lb[2], -3.14159265, -3.14159265, -3.14159265)
54 ub = (aabb_ub[0], aabb_ub[1], aabb_ub[2], 3.14159265, 3.14159265, 3.14159265)
55 return rng.uniform(lb, ub).tolist()
56
57
58@dataclass
59class PlaceSampler(Sampler):
60 curr_state: RavenState
61 env: RavenYCBEnv
62 object_name_to_place_on_or_inside: str
63
64 def sample(self, rng: np.random.Generator) -> List[float]:
65 """Simply return some value within the AABB of the object, and at any
66 orientation."""
67 min_drop_height = 0.01
68 max_drop_height = 0.35
69 aabb_lb, aabb_ub = get_aabb_bounds(
70 self.curr_state, self.env,
71 self.object_name_to_ place_on_or_inside
72 )
73 lb = (
74 aabb_lb[0],
75 aabb_lb[1],
76 aabb_lb[2] + min_drop_height,
77 -3.14159265,
78 -3.14159265,
79 -3.14159265,
80 )
81 ub = (
82 aabb_ub[0],
83 aabb_ub[1],
84 aabb_ub[2] + max_drop_height,
85 3.14159265,
86 3.14159265,
87 3.14159265,
88 )
89 return rng.uniform(lb, ub).tolist()
90
91
92@dataclass
93class PourSampler(Sampler):
94 curr_state: RavenState
95 env: RavenYCBEnv
96 holding_obj_name: str
97 obj_name_to_pour_into_or_ontop: str
98
99 def sample(self, rng: np.random.Generator) -> List[float]:
100 """Return a position sampled from above obj_name_to_pour_ into_or_ontop
101 and a pitch randomly sampled between bounds."""
102 # We sample (x, y, z, pitch.)
103 obj_holding_lb, obj_holding_ub = get_aabb_bounds(
104 self.curr_state, self.env, self.holding_obj_name
105 )
106 min_pour_height = obj_holding_ub[2] - obj_holding_lb[2]
107 max_pour_height = min_pour_height * 2
108 aabb_lb, aabb_ub = get_aabb_bounds(
109 self.curr_state, self.env, self.obj_name_to_pour _into_or_ontop
110 )
111 pitch_range = (-3.14, 3.14)
112 lb = (aabb_lb[0], aabb_lb[1], aabb_lb[2] + min_pour_height, pitch_range[0])
113 ub = (aabb_ub[0], aabb_ub[1], aabb_ub[2] + max_pour_height, pitch_range[1])
114 return rng.uniform(lb, ub).tolist()
115
116Your goal is to generate a python function that returns a plan that performs the provided task. This function can
117use helper functions that must be defined within the scope of the function itself.
118
119The main function should be named EXACTLY ‘gen_plan‘, and it should take in only one parameter corresponding to the environment state as input. Do not change the names. Do not create any additional classes or overwrite any existing ones. You are only allowed to create helper functions inside the ‘gen_plan‘ function.
120
121Current provided goal: {provided_task_goal}
122
123Here is an example ‘gen_plan‘ function for a different task:
124#define user
125Init state:
126bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
127banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
128plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
129table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
130Task goal: put the lemon on the plate and the banana on the table
131
132#define assistant
133‘‘‘python
134def gen_plan(initial:RavenState):
135 plan = []
136 place_pose = PlaceSampler(initial, env, "plate").sample(rng)
137 lemon_grasp = GraspSampler(initial, env, "lemon").sample(rng)
138 plan.append(Action("pick", [[lemon], lemon_grasp]))
139 plan.append(Action("place_ontop", [[lemon, plate], place_pose]))
140 place_pose = PlaceSampler(initial, env, "table").sample(rng)
141 banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
142 plan.append(Action("pick", [[banana], banana_grasp]))
143 plan.append(Action("place_ontop", [[banana, table], RavenPose(x=x_place, y=y_place, z=0.02)]))
144 return plan
145‘‘‘
146
147Here is another example for a different goal involving the same objects and initial state. Note that in this case, the bowl is blue.
148Also note that the banana needs to be rotated so that it fits into the bowl; in general you should pay careful
149attention to any angular constraints that might be important for solving different tasks.
150
151#define user
152Init state:
153bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
154banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
155plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
156table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
157Task goal: serve the banana inside the blue thing after drying it by placing on the plate
158
159#define assistant
160‘‘‘python
161def gen_plan(initial:RavenState):
162 plan = []
163 place_pose = PlaceSampler(initial, env, "plate").sample(rng)
164 banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
165 plan.append(Action("pick", [[banana], grasp]))
166 plan.append(Action("place_ontop", [[banana, plate], place_pose]))
167 bowl_bounds = get_aabb_bounds(initial, env, "bowl")
168 place_pose = PlaceSampler(initial, env, "bowl").sample(rng)
169 bowl_center_pose = get_obj_center(initial, env)
170 # Drop the object in the exact center of the bowl
171 place_pose[0] = bowl_center_pose[0]
172 place_pose[1] = bowl_center_pose[1]
173 place_pose[2] = bowl_center_pose + 0.05
174 # Make sure to pick an orientation such that the banana will fit!
175 place_pose.roll = 1.5
176 banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
177 plan.append(Action("pick", [[banana], grasp]))
178 plan.append(Action("place_inside", [[banana, bowl], place_pose]))
179 return plan
180‘‘‘
181
182Here is yet another example for a different task:
183mug: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
184potted_meat_can: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
185plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
186table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
187Task goal: serve spam from its can into the cup
188
189#define assistant
190‘‘‘python
191def gen_plan(initial:RavenState):
192 plan = []
193 spam_grasp = GraspSampler(initial, env, "potted_meat_can").sample(rng)
194 pour_params = PourSampler(initial, env, "potted_meat_can", "cup").sample(rng)
195 pour_params[3] = 2.0 # necessary for the contents to really fall into the cup.
196 plan.append(Action("pick", [[potted_meat_can], spam_grasp]))
197 plan.append(Action("pour", [[potted_meat_can, mug], pour_params]))
198 return plan
199‘‘‘
200
201
202Make sure to enclose your output with ‘‘‘python <output gen_plan(initial:Ravenstate): function> ‘‘‘ (ignore the angle brackets - those are just for illustrative purposes).’